Check the limit of the function f(x,y) = 3x^2y/(x^2 + y^2) at origin exist or not

1
Expert's answer
2022-02-22T08:22:36-0500

Let us use the polar coordinates "\begin{cases} x=r\cos \theta \\ y=r\sin\theta \end{cases}" , then the limit "(x,y)\to (0,0)" corresponds to the limit "r\to 0". We have

"f(r,\theta)=\frac{3r^3\cos^2\theta \sin \theta}{r^2}=3r\cos^2\theta \sin \theta"

We see that the limit "\lim_{r\to 0} f(r,\theta) =0" exists and independent of "\theta" (as "3\cos^2\theta \sin \theta" is a bounded quantity, so "\lim_{r\to 0}r\cdot (\text{bounded quantity})=0")